# Local Causal Discovery for Structural Evidence of Direct Discrimination

### v1

@@ 6 Real-World Case Study: Fairness in Liver Transplant Allocation
Liver transplantation is a critical therapeutic option for patients with end-stage chronic liver disease and acute liver failure. In the US, demand for liver transplants significantly surpasses the supply of available donor livers [55]. Patients who require a liver transplant are placed on a national waiting list managed by the United Network for Organ Sharing (UNOS). [^6] https://unos.org/ Upon availability of a donor liver, the distribution policy sorts waitlisted patients by multiple prioritization criteria, such as medical urgency, compatibility, and location, while balancing competing objectives such as perceived policy fairness [43]. The US liver transplant system has seen several key policy changes designed to optimize distribution and improve patient outcomes [57] (see Appendix F). The model for end-stage liver disease (MELD) is one facet of this policy, used to determine priority based on illness severity [51]. Despite efforts to improve fair allocation [40], it is widely recognized that US organ allocation suffers from disparities on the basis of race, sex, and socioeconomic status [101].
While a thorough assessment of unfairness in liver transplantation is out of scope, we take sex-based disparities as a case study. Over time, women have been less likely than men to receive a transplant [52, 27, 75, 1]. Women are less represented on liver transplant waitlists primarily due to lower chronic liver disease rates [75, 56]. Height differences, MELD exception points, and underestimated MELD scores have been raised as potential explanations for sex-based disparities [1]. Prior investigations measure these disparities via statistical association [53], which cannot disentangle causal mechanisms [39]. To the best of our knowledge, there have been limited results on sex-based disparities using US national data collected after March 2018 [20]. To address the lack of causal analysis, we apply LD3 to recover structural evidence of direct sex-based discrimination in transplant allocation.
Detecting Discrimination
@@
We utilize the National Standard Transplant Analysis and Research (STAR) dataset [70] containing patient-level information on waitlisted transplant candidates, recipients, and donors from the Organ Procurement & Transplantation Network (OPTN) up to October 2022. We filter to include adult patients awaiting or having undergone a whole liver transplant from deceased donors in 2017-2019 and 2020-2022, keeping features relevant for discrimination discovery (Appendix F). Time frames were selected as stable periods without changes in the UNOS allocation policy. Estimation was performed with double machine learning [15]. [^7] Double machine learning was implemented with econml (https://econml.azurewebsites.net) using random forest classifiers with a 70%/30% train-test split (https://scikit-learn.org). We assumed that no confounders for $\{\mathbf{M},Y\}$ that are descended from the exposure open backdoor paths when conditioned on [66, 62, 35]. Under all experimental settings and time frames, results from LD3 suggest that the effects of sex on receiving a liver are not fully explained by observed confounding and mediation (Table 2). Given the observed variables, causal evidence of direct discrimination is detected (SDC = 1) and corroborated by non-zero weighted CDE estimates with significant $p$-values. In all settings, $\mathbf{A}_{\text{{DE}}}$ contains key factors used in the liver distribution policy, including initial MELD, initial age, region, and active exception case [70]. Consistent results validate the reliability of LD3 in this complex domain. Our results are concordant with prior observations that differences in body size and medical condition contribute to sex-based disparities [54], as weight and diagnosis are in $\mathbf{A}_{\text{{DE}}}$ in all settings. We also highlight that predicted $\mathbf{A}_{\text{{DE}}}$ indicate potential discrimination with respect to ethnicity, education, and payment method. These socioeconomic factors may warrant further investigation for future policy interventions. These preliminary results are only qualitative indicators that further fairness analyses should take place.
Unos Policy (2017–2019) Unos Policy (2020–2022) LD3 ($\alpha=0.01$) LD3 ($\alpha=0.05$) LD3 ($\alpha=0.01$) LD3 ($\alpha=0.05$) $\mathbf{A}_{\text{{DE}}}$ active exception case, diagnosis, education, ethnicity, exception type, initial MELD, initial age, region, weight active exception case, diagnosis, education, ethnicity, exception type, initial MELD, initial age, recipient blood type, region, weight active exception case, diagnosis, education, ethnicity, initial MELD, initial age, payment method, region, weight active exception case, diagnosis, education, ethnicity, initial MELD, initial age, payment method, region, weight SDC 1 1 1 1 WCDE 0.033 [0.016,0.049] 0.025 [0.008,0.042] 0.032 [0.018,0.047] 0.032 [0.018,0.047] $p$-value 0.000 0.003 0.000 0.000
Table 2: Results obtained with nonparametric $\chi^{2}$ independence tests and double machine learning. Total features is 18 (Appendix F). For time period 2017-2019, total observations is $n=21101$ patients. For time period 2020-2022, total observations is $n=22807$ patients. With no covariate control, SDC = 1 and the weighted CDE is 0.049 [0.033,0.066] ($p$-value = 0.000).
@@
SDC and weighted CDE results for PC [38] and LDECC [31] are reported in Tables E.5 and E.6. PC and LDECC qualitatively agree with LD3 on both discrimination metrics, though at times with higher weighted CDE values. However, PC, LDECC, and LD3 had relatively low agreement in terms of $\mathbf{A}_{\text{{DE}}}$. Adjustment set cardinality was lower for PC and LDECC, and was zero for PC in 2017-2019 ($\alpha=0.05$) and one for LDECC the same period ($\alpha=0.01$). Most settings for PC and LDECC returned adjustment sets that omitted key expected variables, such as known policy criteria (e.g., initial MELD score, active exception case, and initial age). PC results for both significance levels show multiple untrustworthy directed edges (e.g., weight $\to$ sex, height $\to$ sex, blood type $\to$ age, education $\to$ height, body mass index $\to$ height), undermining the credibility of results. Likewise, LDECC predicts that the outcome has children, which is known to be untrue.
Efficiency Gains with LD3
Discovery with sequential independence tests is infamously slow. Relative to LD3, vanilla PC performed 479–1021$\times$ more tests and took 2295–5870$\times$ longer to execute on an Apple M2 Pro Chip (Table E.7). LDECC performed 42–984$\times$ more tests and took 197–5774$\times$ longer to run. Speedups of this magnitude will be impactful for researchers who require relatively slow nonparametric tests with large sample sizes (e.g., [28, 29, 99, 71]) and for high-dimensional feature spaces, especially when access to compute is limited. Speedups reported throughout this work meet or exceed those reported in limited work on hardware acceleration for embarrassingly parallel constraint-based discovery, comparing single-threaded CPUs, multi-threaded CPUs, and GPUs [97, 32, 10]. For nonparametric discovery to increase its real-world impact, it will need to efficiently handle increasingly large data within reasonable academic compute limits.
@@ 7 Conclusion
This work advocates for increased practicality in causal fairness pipelines, with a mind to computational efficiency and implementability. We propose a time-efficient and asymptotically correct local discovery method for identifying two qualitative measures of direct discrimination: the structural direct criterion and the weighted CDE. The number of independence tests performed is worst-case linear in the number of nodes, while comparable baselines are worst-case quadratic and exponential. We illustrate the utility of our approach on a real-world clinical fairness problem, providing causal structural evidence of direct sex-based discrimination in the national liver transplant allocation policy.
@@
Future work could extend LD3 to allow for $Y$ with descendants. The assumption that all parents of $Y$ are observed is sufficient but not necessary for CDE identification and could be replaced with a different criterion. The performance of LD3 will be constrained by the properties of the selected independence test, and any future improvements in statistical independence testing will be of great benefit to the causal discovery community. Future work could consider discovery for other fairness estimands, such as the natural effects [62] and path-specific effects [3].
Broader Impacts and Ethical Implications
Matters of unfairness and direct discrimination have real, profound impacts on human well-being and are often challenging to detect in real-world retrospective data. Researchers who deploy automated methods like LD3 should be prudent in interpreting and publicly communicating the statistical significance and uncertainty of their findings, whether evidence for or against discrimination is found.
@@
Appendix A Supplemental Figures and Tables
Disparate Treatment Disparate Impacts
Direct Causal Mechanism Indirect Causal Mechanism Spurious Association
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}X}$$Y$$Z$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}X}$$Y$$Z$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}X}$$Y$$Z$
$X\sim\text{Bern}(0.3)$ $X\sim\text{Bern}(0.3)$ $X\sim\text{Bern}(0.3+0.1Z)$
$Z\sim\text{Bern}(0.2+0.3{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}X})$ $Z\sim\text{Bern}(0.2+0.3{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}X})$ $Z\sim\text{Bern}(0.2)$
$Y\sim\text{Bern}(0.1+0.2{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}X})$ $Y\sim\text{Bern}(0.1+0.2Z)$ $Y\sim\text{Bern}(0.1+0.2Z)$
Table A.1: Example data generating processes under direct, indirect, and spurious structural mechanisms of unfairness. Adapted from Bareinboim & Plecko [4]. Let $X$ denote a protected attribute, $Y$ an outcome of interest, and $Z$ a third measured variable.
Appendix B Further Discussion on Direct Effect Estimation
B.1 Comparing the Weighted Controlled Direct Effect to the Natural Direct Effect
We provide additional discussion on the weighted CDE and its most popular counterpart, the NDE. Recall that the CDE and weighted CDE are defined as follows:
$CDE(\mathbf{m})$ $\coloneqq\mathbb{E}[Y_{x,\mathbf{m}}-Y_{x^{*},\mathbf{m}}]$ (5)
$=\mathbb{E}[Y|do(X=x,\mathbf{M}=\mathbf{m})]-\mathbb{E}[Y|do(X=x^{*},\mathbf{M}=\mathbf{m})]$ (6)
$=\sum_{\mathbf{s}}\big{[}\mathbb{E}[Y|x,\mathbf{s},\mathbf{m}]-\mathbb{E}[Y|x^{*},\mathbf{s},\mathbf{m}]\big{]}P(\mathbf{s}).$ (7)
$WCDE$ $\coloneqq\sum_{\mathbf{m}}\sum_{\mathbf{s}}\big{[}\big{[}\mathbb{E}[Y|x,\mathbf{s},\mathbf{m}]-\mathbb{E}[Y|x^{*},\mathbf{s},\mathbf{m}]\big{]}P(\mathbf{s})\big{]}P(\mathbf{m}).$ (8)
Definition B.1 (Average NDE, Pearl [62]).
@@ B.1 Comparing the Weighted Controlled Direct Effect to the Natural Direct Effect
The NDE measures the expected change in outcome as the exposure changes, with mediators fixed at the values that they would have been per individual prior to this change in exposure. The NDE captures the portion of the total causal effect that would be transmitted to the outcome if the mediators did not respond to changes in the exposure.
For values of the exposure $x$ and $x^{*}$, values $\mathbf{s}$ of covariates $\mathbf{S}$, and values $\mathbf{m}$ of mediators $\mathbf{M}$, the average NDE is
$NDE(x,x^{*};Y)$ $=\mathbb{E}[Y_{x,\mathbf{M}_{x}^{*}}-Y_{x^{*},\mathbf{M}_{x}^{*}}].$ (9)
The average NDE in Markovian models is identifiable from observational data per Corollary 2 in Pearl [59], allowing counterfactual expressions to be replaced by probabilistic estimands:
$NDE(x,x^{*};Y)$ $=\sum_{\mathbf{s}}\sum_{\mathbf{m}}\big{[}\big{[}\mathbb{E}[Y|x,\mathbf{m}]-\mathbb{E}[Y|x^{*},\mathbf{m}]\big{]}P(\mathbf{m}|x^{*},\mathbf{s})\big{]}P(\mathbf{s}).$ (10)
@@
We note that the weighted CDE and NDE differ only by the way that we weigh the CDE: for the weighted CDE, we weigh by $P(\mathbf{m})$, and for the NDE, we weigh by $P(\mathbf{m}|x^{*},s)$. This stems from the fact that the CDE is a function of interventional distributions, while the NDE is a counterfactual quantity. The NDE requires probabilities of nested counterfactuals and a cross-world counterfactual assumption [82, 60, 62, 2]. Nevertheless, Pearl [59] showed that the NDE can be expressed with do-notation and reduced to a “do-free” expression by valid covariate adjustment, allowing for identification and estimation [60]. Unlike the NDE, the CDE does not admit a valid decomposition of the total causal effect into direct and indirect effects for the general nonparametric setting [59]. Conversely, the total effect is always equal to the difference of the NDE and the reverse transition of the natural indirect effect (NIE), even in the presence of nonlinearities and exposure-mediator interactions [62]. However, in the absence of exposure-mediator interaction, the CDE and NDE are equivalent and the CDE can be used for total effect decomposition [89]. The guaranteed decomposability of the total effect into the NDE and NIE is a primary reason that the natural effects are employed in causal fairness analysis [69]. However, when only evidence of direct discrimination is needed, it suffices to estimate the CDE [98]. In practice, the CDE has traditionally been in more frequent use in policy evaluation settings than the NDE [89].
The present work is centrally concerned with direct discrimination only. For our current objectives, we favor the weighted CDE as a metric because it requires fewer assumptions over the data generating process and requires less prior structural knowledge than the NDE [89]. For example, the NDE requires that no confounder for $\{\mathbf{M},Y\}$ is descended from $X$ [62], as the same covariates that deconfound $\{\mathbf{M},Y\}$ must also be used to identify the counterfactual quantity $P(\mathbf{m}|x^{*},s)$. As explained in Appendix C of Pearl [62], if this covariate set contains a variable that is descended from the exposure, $P(\mathbf{m}|x^{*},s)$ is not identifiable even under random treatment assignment. While adjusting for confounders for $\{\mathbf{M},Y\}$ that are descended from the exposure can induce bias in CDE estimation, this bias is addressable when given an appropriate estimator [66, 62, 35]. Therefore, the presence or absence of such variables is something that can be assumed at the estimation stage at the discretion of the user, but does not impact the behavior of LD3. The absence of confounders of this form is challenging to verify [89], and may be unrealistic in real-world data. This limits the range of structures for which NDE estimates can be obtained. Additionally, the identifiability of the NDE requires that $\mathbf{S}$ also deconfounds $\{X,\mathbf{M}\}$, adding an extra burden of graphical knowledge. See Andrews & Didelez [2] for a discussion of the cross-world counterfactual assumption in NDE identification. Though the NDE has useful properties for certain forms of analysis, we restrict our attention to the weighted CDE in the present work for these reasons.
Appendix C Proofs
C.1 Correctness of Algorithm 1

### v3

@@ 6.2 Sex and Liver Transplant Allocation
We apply LD3 to a real-world case study in the US healthcare system: fairness in liver transplant allocation. Liver transplantation is a critical therapeutic option for patients with end-stage chronic liver disease and acute liver failure. Demand significantly surpasses supply for donor livers (SRTR 2020), and patients are placed on a national waiting list managed by the United Network for Organ Sharing (UNOS). The distribution policy sorts waitlisted patients by multiple criteria, such as medical urgency, compatibility, and location (Latt, Niazi, and Pyrsopoulos 2022). Several key policy changes have sought to optimize distribution and improve patient outcomes (Papalexopoulos et al. 2023; see Appendix E.2), including the model for end-stage liver disease (MELD; Malinchoc et al. 2000). Despite efforts to increase fairness (Kim et al. 2022), it is widely recognized that US organ allocation suffers from disparities (Zhang et al. 2018).
We explored potential sex-based discrimination in liver allocation. Sex-based disparities have been observed as statistical associations (Mathur et al. 2011; Gordon and Ladner 2012; Sarkar et al. 2015; Oloruntoba and Moylan 2015; Allen et al. 2018; Nephew and Serper 2021), but have not been explored through a causal lens. We use the National Standard Transplant Analysis and Research (STAR) dataset (OPTN 2024) for adult patients during 2017-2019 ($n=21,101$) and 2020-2022 ($n=22,807$). See Appendix E.2 for feature selection and summary statistics.
Results
@@
Under all experimental settings and time frames, results from LD3 suggest that the effects of sex on receiving a liver were not fully explained by observed confounding and mediation (Figure 5, Table E.2). Causal evidence of direct discrimination was detected (SDC = 1) and corroborated by non-zero WCDE with significant $p$-values ($<0.005$). In all settings, $\mathbf{A}_{\text{{DE}}}$ contains key factors used in the liver distribution policy, including initial MELD, initial age, region, and active exception case (OPTN 2024). Our results are concordant with prior observations that differences in body size and medical condition contribute to sex-based disparities (Nephew et al. 2017), as weight and diagnosis are in $\mathbf{A}_{\text{{DE}}}$ in all settings. Predicted $\mathbf{A}_{\text{{DE}}}$ also indicate potential discrimination with respect to ethnicity, education, and payment method. These socioeconomic factors may warrant further investigation for future policy interventions.
@@
PC and LDECC qualitatively agree with LD3 on both discrimination metrics (Figure 5, Tables E.3, E.4). However, PC, LDECC, and LD3 had relatively low agreement in terms of $\mathbf{A}_{\text{{DE}}}$. Adjustment set cardinality was lower (and at times zero) for PC and LDECC. Most settings for PC and LDECC returned $\mathbf{A}_{\text{{DE}}}$ that omitted key expected variables, such as known policy criteria (e.g., initial MELD score, active exception case, and initial age). At both significance levels, PC returned multiple untrustworthy edges (e.g., weight $\to$ sex, height $\to$ sex, blood type $\to$ age, education $\to$ height, body mass index $\to$ height), undermining the credibility of results. Likewise, LDECC predicts that the outcome has children, which is known to be untrue. PC performed 479–1021$\times$ more tests and took 2295–5870$\times$ longer to execute than LD3 (Table E.5). LDECC performed 42–984$\times$ more tests and took 197–5774$\times$ longer to run.
@@ 7 Conclusion
This work advocates for increased practicality in causal fairness pipelines. We propose a time-efficient and asymptotically correct local discovery method for identifying two qualitative indicators of direct discrimination: the SDC and WCDE. For two real-world causal fairness analyses, LD3 returned more stable and plausible predictions with signficantly better computational efficiency relative to baselines.
@@
Future work could extend LD3 to allow for $Y$ with descendants. The assumption that all parents of $Y$ are observed is sufficient but not necessary for CDE identification and could be replaced with a different criterion. LD3 cannot differentiate $\mathbf{Z}_{1}$ from $\mathbf{Z}_{3}$, which would enable analysis of indirect and spurious discrimination. Future work could consider discovery for other fairness estimands, such as the natural effects (Pearl 2014) and path-specific effects (Avin, Shpitser, and Pearl 2005).
Acknowledgments
This research was supported by NSF 2212175; NIH RF1AG084178, R01AG076448, R01AG080624, R01AG076234, R01AG080991 and RF1AG072449; and the NSF Graduate Research Fellowship (J. Maasch).
@@ Acknowledgments
Appendix A Further Discussion on Direct Effect Estimation
A.1 Comparison of Direct Effect Measures
Here we explore the differences and similarities between the CDE and alternative direct effect measures, and provide justification for our use of the CDE in this work.
Alternative Direct Effect Measures
We elaborate on the comparison between the CDE and its most popular counterpart, the NDE (Pearl 2001). [^11] The Ctf-DE is an extension of the NDE; see Zhang and Bareinboim (2018). The CDE measures the expected change in outcome as the exposure changes when mediators $\mathbf{M}$ are uniformly fixed to $\mathbf{m}$, a constant value over the population (as simulated by conditioning on $\mathbf{M}$; Pearl 2012a). We express the CDE in $do$-notation as
$\mathrm{CDE}(\mathbf{m})$ $\coloneqq\mathbb{E}[Y_{x,\mathbf{m}}-Y_{x^{*},\mathbf{m}}]$ (7)
$=\mathbb{E}[Y|do(X=x,\mathbf{M}=\mathbf{m})]-\mathbb{E}[Y|do(X=x^{*},\mathbf{M}=\mathbf{m})].$ (8)
Assuming access to a covariate set $\mathbf{S}$ that simultaneously controls for confounding of the exposure-outcome and mediator-outcome relationships such that the conditions described in Definition 5 are satisfied, [^12] When a single set $\mathbf{S}$ is not sufficient for valid control of confounding for both $\{X,Y\}$ and $\{\mathbf{M},Y\}$, we can apply a more complex formula that requires two covariate sets. See Vanderweele 2011 (Proposition 1) and Pearl 2014 (Appendix C) for elaboration. the CDE is identifiable from observational data as
$\sum_{\mathbf{s}}\big{(}\mathbb{E}[Y|x,\mathbf{s},\mathbf{m}]-\mathbb{E}[Y|x^{*},\mathbf{s},\mathbf{m}]\big{)}P(\mathbf{s}).$ (9)
To obtain a unique estimate irrespective of potential interactions between $X$ and $\mathbf{M}$, we marginalize over the mediator values:
$\mathrm{WCDE}$ $\coloneqq\sum_{\mathbf{m}}\sum_{\mathbf{s}}\big{(}\mathbb{E}[Y|x,\mathbf{s},\mathbf{m}]-\mathbb{E}[Y|x^{*},\mathbf{s},\mathbf{m}]\big{)}P(\mathbf{s})P(\mathbf{m}).$ (10)
@@ A.1 Comparison of Direct Effect Measures
In contrast, the NDE measures the expected change in outcome as the exposure changes, with mediators fixed at the values that they would have been per individual prior to this change in exposure. The NDE captures the portion of the total causal effect that would be transmitted to the outcome if the mediators did not respond to changes in the exposure.
Definition A.1 (Natural direct effect (NDE), Pearl 2014).
We express the average NDE as
$\mathrm{NDE}(x,x^{*};Y)$ $=\mathbb{E}[Y_{x,\mathbf{M}_{x}^{*}}-Y_{x^{*},\mathbf{M}_{x}^{*}}].$ (11)
where $\mathbf{M}_{x}^{*}$ takes the value that the individual would have attained under $X=x^{*}$. The NDE requires probabilities of nested counterfactuals and a cross-world counterfactual assumption (Shpitser and VanderWeele 2011; Pearl 2012b, 2014; Andrews and Didelez 2021). Nevertheless, Pearl (2001) showed that the NDE can be expressed with do-notation and reduced to a do-free expression by valid covariate adjustment, allowing for identification and estimation (Pearl 2012b). The average NDE in Markovian models is identifiable from observational data per Corollary 2 in Pearl (2001), allowing counterfactual expressions to be replaced by probabilistic estimands:
$\mathrm{NDE}(x,x^{*};Y)$ $=\sum_{\mathbf{m}}\sum_{\mathbf{s}}\big{(}\mathbb{E}[Y|x,\mathbf{s},\mathbf{m}]-\mathbb{E}[Y|x^{*},\mathbf{s},\mathbf{m}]\big{)}P(\mathbf{s})P(\mathbf{m}|x^{*},\mathbf{s}).$ (12)
Note that the WCDE and NDE differ only by the way that we weigh the CDE: for the WCDE, we weigh by $P(\mathbf{m})$, and for the NDE, we weigh by $P(\mathbf{m}|x^{*},s)$. This stems from the fact that the CDE is a function of interventional distributions, while the NDE is a counterfactual quantity. Meanwhile, the Ctf-DE measures NDEs conditioned on the subpopulation $X=x$ (Zhang and Bareinboim 2018).
Total Effect Decomposition
@@
Unlike the NDE, the CDE does not admit a valid decomposition of the total causal effect into direct and indirect effects for the general nonparametric setting (Pearl 2001). Conversely, the total effect is always equal to the difference of the NDE and the reverse transition of the natural indirect effect (NIE), even in the presence of nonlinearities and exposure-mediator interactions (Pearl 2014). However, in the absence of exposure-mediator interaction, the CDE and NDE are equivalent and the CDE can be used for total effect decomposition (Vanderweele 2011). Guaranteed total effect decomposability is a primary reason that the natural effects are employed in CFA (Plečko and Bareinboim 2024). However, when qualitative indicators of direct discrimination are all that is needed, it suffices to estimate the CDE (Zhang and Bareinboim 2018).
A.2 Graphical Criteria for Direct Effect Identification
Sufficient and necessary conditions for identification vary by estimand, but generally require covariate adjustment (VanderWeele 2019). Formal graphical rules can allow the VAS to be read directly from causal structures, even in the presence of latent confounding (Perkovic et al. 2015; Maathuis and Colombo 2015; Runge 2021). Graphical criteria for valid adjustment have been proposed for diverse causal quantities (e.g., Pearl’s backdoor criterion for total effect estimation; Pearl 1995). Often, multiple valid criteria exist per estimand. These vary in the amount of prior structural knowledge required, and criteria that require less knowledge are easier to implement in practice (VanderWeele and Shpitser 2013; Guo, Lundborg, and Zhao 2022).
Several graphical criteria have been proposed for the identification of direct effects and related quantities. Shpitser and Pearl (2006) present a complete graphical criterion for conditional interventional distributions. Pearl (2001) defines a graphical criterion for identifying the NDE from observational data (Corollary 1). Discussion, debate, and confusion around the sufficient and necessary conditions for NDE identification have inspired a flurry of papers (Imai, Keele, and Yamamoto 2010; Imai, Keele, and Tingley 2010), as summarized in Pearl (2014). Shpitser and VanderWeele (2011) present a complete graphical criterion for the NDE that reconciles some debate, showing where common criteria coincide.
Graphical criteria for direct effect estimation are often challenging for practitioners to implement (Shpitser and VanderWeele 2011). To increase the practicality of covariate adjustment in complex and low-knowledge domains, we advocate for the development of graphical criteria that are (1) relatively intuitive and practical to implement in applied research, and (2) are satisfied by knowledge that can be reliably learned through data-driven methods. Thus, we apply the causal partition taxonomy of Maasch et al. (2024) to provide local graphical intuition for fast and practical adjustment set discovery.
Appendix B Proofs
B.1 Proof of Theorem 1 (Correctness of Algorithm 1)